# Weyl distributions, spectral properties, and circulant approximation results for quaternion block multilevel Toeplitz matrix sequences

### v1

@@ 2.4 Function theory preliminaries and notation
When dealing with spectral distributions of matrix sequences, we use the following set.
@@ 4.5 Spectral localization of Hermitian matrices via the $2s\times 2s$ complex symbols
$Z(\bm{\theta}):=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(p_{\mathrm{R}}(\bm{\rho})^{(0)}+p_{\mathrm{R}}(\bm{\rho})^{(1)}\mathbf{i}\bigr)e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle},\qquad W(\bm{\theta}):=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(p_{\mathrm{R}}(\bm{\rho})^{(2)}+p_{\mathrm{R}}(\bm{\rho})^{(3)}\mathbf{i}\bigr)e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle}.$ (4.25)
For our aims, we need the symplectic embedding in block form $\Phi_{1}:\mathbb{H}^{s\times t}\to\mathbb{C}^{(2s)\times(2t)}$.
@@
$Z(\bm{\theta}):=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(p_{\mathrm{R}}(\bm{\rho})^{(0)}+p_{\mathrm{R}}(\bm{\rho})^{(1)}\mathbf{i}\bigr)e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle},\qquad W(\bm{\theta}):=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(p_{\mathrm{R}}(\bm{\rho})^{(2)}+p_{\mathrm{R}}(\bm{\rho})^{(3)}\mathbf{i}\bigr)e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle},$
@@
$\Pi_{L}\,U_{L}\,C_{\bm{n}}(p_{\mathrm{R}})\,U_{R}^{*}\,\Pi_{R}^{*}=\bigoplus_{\bm{k}\in\mathrm{Fix}}\bigl(Z(\bm{\theta}_{\bm{k}})+W(\bm{\theta}_{\bm{k}})\mathbf{j}\bigr)\ \ \oplus\ \bigoplus_{\{\bm{k},-\bm{k}\}\subset\Lambda_{\bm{n}}\setminus\mathrm{Fix}}\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(\bm{\theta}_{\bm{k}})\mathbf{j}\\[2.0pt] W(\bm{\theta}_{-\bm{k}})\mathbf{j}&Z(\bm{\theta}_{-\bm{k}})\end{bmatrix}$, (4.26)
@@
$\Phi\!\left(\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(\bm{\theta}_{\bm{k}})\mathbf{j}\\[2.0pt] W(\bm{\theta}_{-\bm{k}})\mathbf{j}&Z(\bm{\theta}_{-\bm{k}})\end{bmatrix}\right)=\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&0&0&W(\bm{\theta}_{\bm{k}})\\[2.0pt] 0&\overline{Z(\bm{\theta}_{\bm{k}})}&-\overline{W(\bm{\theta}_{\bm{k}})}&0\\[2.0pt] 0&W(\bm{\theta}_{-\bm{k}})&Z(\bm{\theta}_{-\bm{k}})&0\\[2.0pt] -\overline{W(\bm{\theta}_{-\bm{k}})}&0&0&\overline{Z(\bm{\theta}_{-\bm{k}})}\end{bmatrix}$, (4.27)
@@
$\Phi\bigl(Z(\bm{\theta}_{\bm{k}})+W(\bm{\theta}_{\bm{k}})\mathbf{j}\bigr)=\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(\bm{\theta}_{\bm{k}})\\[2.0pt] -\overline{W(\bm{\theta}_{\bm{k}})}&\overline{Z(\bm{\theta}_{\bm{k}})}\end{bmatrix}$. (4.28)
@@
$Z(\bm{\theta}_{\bm{k}})=\widehat{S}_{0}(\bm{k})+\widehat{S}_{1}(\bm{k})\,\mathbf{i},\qquad W(\bm{\theta}_{\bm{k}})=\widehat{S}_{2}(\bm{k})+\widehat{S}_{3}(\bm{k})\,\mathbf{i}.$
Thus the fixed blocks are indeed $Z(\bm{\theta}_{\bm{k}})+W(\bm{\theta}_{\bm{k}})\mathbf{j}$ and the paired blocks are
$\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(\bm{\theta}_{\bm{k}})\mathbf{j}\\[2.0pt] W(\bm{\theta}_{-\bm{k}})\mathbf{j}&Z(\bm{\theta}_{-\bm{k}})\end{bmatrix}$,
@@
$\Phi\bigl(Z(\bm{\theta}_{\bm{k}})+W(\bm{\theta}_{\bm{k}})\mathbf{j}\bigr)=\Phi\bigl(Z(\bm{\theta}_{\bm{k}})\bigr)+\Phi\bigl(W(\bm{\theta}_{\bm{k}})\mathbf{j}\bigr)=\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(\bm{\theta}_{\bm{k}})\\[2.0pt] -\overline{W(\bm{\theta}_{\bm{k}})}&\overline{Z(\bm{\theta}_{\bm{k}})}\end{bmatrix}$,
@@
$\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(\bm{\theta}_{\bm{k}})\mathbf{j}\\[2.0pt] W(\bm{\theta}_{-\bm{k}})\mathbf{j}&Z(\bm{\theta}_{-\bm{k}})\end{bmatrix}=\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&0\\[2.0pt] 0&Z(\bm{\theta}_{-\bm{k}})\end{bmatrix}\;+\;\begin{bmatrix}0&W(\bm{\theta}_{\bm{k}})\mathbf{j}\\[2.0pt] W(\bm{\theta}_{-\bm{k}})\mathbf{j}&0\end{bmatrix}$.
@@
$Z(\bm{\theta})=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(p_{\mathrm{R}}(\bm{\rho})^{(0)}+p_{\mathrm{R}}(\bm{\rho})^{(1)}\mathbf{i}\bigr)e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle},\qquad W(\bm{\theta})=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(p_{\mathrm{R}}(\bm{\rho})^{(2)}+p_{\mathrm{R}}(\bm{\rho})^{(3)}\mathbf{i}\bigr)e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle},$
$G_{\mathrm{R}}\bigl(p_{\mathrm{R}}(\bm{\theta})\bigr):=\begin{bmatrix}Z(\bm{\theta})&W(\bm{\theta})\\[2.0pt] -\overline{W(-\bm{\theta})}&\overline{Z(-\bm{\theta})}\end{bmatrix}$.
@@
$\bigl(Z(\bm{\theta}_{\bm{k}})+W(\bm{\theta}_{\bm{k}})\mathbf{j}\bigr)\quad(\bm{k}\in\mathrm{Fix}),\qquad\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(\bm{\theta}_{\bm{k}})\mathbf{j}\\[2.0pt] W(\bm{\theta}_{-\bm{k}})\mathbf{j}&Z(\bm{\theta}_{-\bm{k}})\end{bmatrix}\quad(\{\bm{k},-\bm{k}\}),$
@@
$\Phi\bigl(Z(\bm{\theta}_{\bm{k}})+W(\bm{\theta}_{\bm{k}})\mathbf{j}\bigr)=\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(\bm{\theta}_{\bm{k}})\\[2.0pt] -\overline{W(\bm{\theta}_{\bm{k}})}&\overline{Z(\bm{\theta}_{\bm{k}})}\end{bmatrix}=G_{\mathrm{R}}\bigl(p_{\mathrm{R}}(\bm{\theta}_{\bm{k}})\bigr).$
@@ Acknowledgements
Valerio Loi and Stefano Serra-Capizzano are members of the research group GNCS (Gruppo Nazionale per il Calcolo Scientifico) of INdAM (Istituto Nazionale di Alta Matematica). Furthermore, the third author has been supported by the Theory, Economics and Systems Laboratory (TESLAB) of the Department of Computer Science of the Athens University of Economics and Business and by the PRIN-PNRR project MATHPROCULT (MATHematical tools for predictive maintenance and PROtection of CULTural heritage, Code P20228HZWR, CUP J53D23003780006).

### v2

@@ 4.5 Spectral localization of Hermitian matrices via the $2s\times 2s$ complex symbols
$Z(\bm{\theta}):=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(p_{\mathrm{R}}(\bm{\rho})^{(0)}+p_{\mathrm{R}}(\bm{\rho})^{(1)}\mathbf{i}\bigr)e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle},\qquad W(-\bm{\theta}):=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(p_{\mathrm{R}}(\bm{\rho})^{(2)}+p_{\mathrm{R}}(\bm{\rho})^{(3)}\mathbf{i}\bigr)e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle}.$ (4.25)
Note that the choice of the sign is $W(-\bm{\theta})$ is made to follow the convention of the Fourier Coefficients sign as in Definition 2.14 and Lemma 2.15. For our aims, we need the symplectic embedding in block form $\Phi_{1}:\mathbb{H}^{s\times t}\to\mathbb{C}^{(2s)\times(2t)}$.
@@
$Z(\bm{\theta}):=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(p_{\mathrm{R}}(\bm{\rho})^{(0)}+p_{\mathrm{R}}(\bm{\rho})^{(1)}\mathbf{i}\bigr)e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle},\qquad W(-\bm{\theta}):=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(p_{\mathrm{R}}(\bm{\rho})^{(2)}+p_{\mathrm{R}}(\bm{\rho})^{(3)}\mathbf{i}\bigr)e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle},$
@@
$\Pi_{L}\,U_{L}\,C_{\bm{n}}(p_{\mathrm{R}})\,U_{R}^{*}\,\Pi_{R}^{*}=\bigoplus_{\bm{k}\in\mathrm{Fix}}\bigl(Z(\bm{\theta}_{\bm{k}})+W(-\bm{\theta}_{\bm{k}})\mathbf{j}\bigr)\ \ \oplus\ \bigoplus_{\{\bm{k},-\bm{k}\}\subset\Lambda_{\bm{n}}\setminus\mathrm{Fix}}\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(-\bm{\theta}_{\bm{k}})\mathbf{j}\\[2.0pt] W(-\bm{\theta}_{-\bm{k}})\mathbf{j}&Z(\bm{\theta}_{-\bm{k}})\end{bmatrix}$, (4.26)
@@
$\Phi\!\left(\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(-\bm{\theta}_{\bm{k}})\mathbf{j}\\[2.0pt] W(-\bm{\theta}_{-\bm{k}})\mathbf{j}&Z(\bm{\theta}_{-\bm{k}})\end{bmatrix}\right)=\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&0&0&W(-\bm{\theta}_{\bm{k}})\\[2.0pt] 0&\overline{Z(\bm{\theta}_{\bm{k}})}&-\overline{W(-\bm{\theta}_{\bm{k}})}&0\\[2.0pt] 0&W(-\bm{\theta}_{-\bm{k}})&Z(\bm{\theta}_{-\bm{k}})&0\\[2.0pt] -\overline{W(-\bm{\theta}_{-\bm{k}})}&0&0&\overline{Z(\bm{\theta}_{-\bm{k}})}\end{bmatrix}$, (4.27)
@@
$\Phi\bigl(Z(\bm{\theta}_{\bm{k}})+W(-\bm{\theta}_{\bm{k}})\mathbf{j}\bigr)=\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(-\bm{\theta}_{\bm{k}})\\[2.0pt] -\overline{W(-\bm{\theta}_{\bm{k}})}&\overline{Z(\bm{\theta}_{\bm{k}})}\end{bmatrix}$. (4.28)
@@
$Z(\bm{\theta}_{\bm{k}})=\widehat{S}_{0}(\bm{k})+\widehat{S}_{1}(\bm{k})\,\mathbf{i},\qquad W(-\bm{\theta}_{\bm{k}})=\widehat{S}_{2}(\bm{k})+\widehat{S}_{3}(\bm{k})\,\mathbf{i}.$
Thus the fixed blocks are indeed $Z(\bm{\theta}_{\bm{k}})+W(-\bm{\theta}_{\bm{k}})\mathbf{j}$ and the paired blocks are
$\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(-\bm{\theta}_{\bm{k}})\mathbf{j}\\[2.0pt] W(-\bm{\theta}_{-\bm{k}})\mathbf{j}&Z(\bm{\theta}_{-\bm{k}})\end{bmatrix}$,
@@
$\Phi\bigl(Z(\bm{\theta}_{\bm{k}})+W(\bm{\theta}_{\bm{k}})\mathbf{j}\bigr)=\Phi\bigl(Z(\bm{\theta}_{\bm{k}})\bigr)+\Phi\bigl(W(-\bm{\theta}_{\bm{k}})\mathbf{j}\bigr)=\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(-\bm{\theta}_{\bm{k}})\\[2.0pt] -\overline{W(-\bm{\theta}_{\bm{k}})}&\overline{Z(\bm{\theta}_{\bm{k}})}\end{bmatrix}$,
@@
$\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(-\bm{\theta}_{\bm{k}})\mathbf{j}\\[2.0pt] W(-\bm{\theta}_{-\bm{k}})\mathbf{j}&Z(\bm{\theta}_{-\bm{k}})\end{bmatrix}=\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&0\\[2.0pt] 0&Z(\bm{\theta}_{-\bm{k}})\end{bmatrix}\;+\;\begin{bmatrix}0&W(-\bm{\theta}_{\bm{k}})\mathbf{j}\\[2.0pt] W(-\bm{\theta}_{-\bm{k}})\mathbf{j}&0\end{bmatrix}$.
@@
$Z(\bm{\theta})=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(p_{\mathrm{R}}(\bm{\rho})^{(0)}+p_{\mathrm{R}}(\bm{\rho})^{(1)}\mathbf{i}\bigr)e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle},\qquad W(-\bm{\theta})=\sum_{\bm{\rho}\in\mathcal{F}}\bigl(p_{\mathrm{R}}(\bm{\rho})^{(2)}+p_{\mathrm{R}}(\bm{\rho})^{(3)}\mathbf{i}\bigr)e^{-\mathbf{i}\langle\bm{\rho},\bm{\theta}\rangle},$
$G_{\mathrm{R}}\bigl(p_{\mathrm{R}}(\bm{\theta})\bigr):=\begin{bmatrix}Z(\bm{\theta})&W(-\bm{\theta})\\[2.0pt] -\overline{W(\bm{\theta})}&\overline{Z(-\bm{\theta})}\end{bmatrix}$.
@@
$\bigl(Z(\bm{\theta}_{\bm{k}})+W(-\bm{\theta}_{\bm{k}})\mathbf{j}\bigr)\quad(\bm{k}\in\mathrm{Fix}),\qquad\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(-\bm{\theta}_{\bm{k}})\mathbf{j}\\[2.0pt] W(-\bm{\theta}_{-\bm{k}})\mathbf{j}&Z(\bm{\theta}_{-\bm{k}})\end{bmatrix}\quad(\{\bm{k},-\bm{k}\}),$
@@
$\Phi\bigl(Z(\bm{\theta}_{\bm{k}})+W(-\bm{\theta}_{\bm{k}})\mathbf{j}\bigr)=\begin{bmatrix}Z(\bm{\theta}_{\bm{k}})&W(-\bm{\theta}_{\bm{k}})\\[2.0pt] -\overline{W(-\bm{\theta}_{\bm{k}})}&\overline{Z(\bm{\theta}_{\bm{k}})}\end{bmatrix}=G_{\mathrm{R}}\bigl(p_{\mathrm{R}}(\bm{\theta}_{\bm{k}})\bigr).$
@@ Acknowledgements
Valerio Loi and Stefano Serra-Capizzano are members of the research group GNCS (Gruppo Nazionale per il Calcolo Scientifico) of INdAM (Istituto Nazionale di Alta Matematica).